Research has shown that 55% of new Small Medium Enterprises (SMEs) are started by graduates while 45% are started by non-graduates. It is also known that 70% of SMEs started by graduates are successful i.e they suvived beyond 3 years, while only 10% of those started by non-graduates are successful.


a. What is the probability that a new SME is successfull

b. What is the probability that a new SME is successful and it was not started by a graduate

c. If it is know that a new SME is successful, what is the probability that it was not started by a graduatre


1
Expert's answer
2022-05-17T16:13:39-0400

Let "A" denote the event "SME is started by graduate".

Let "B" denote the event "SME is started by non-graduate".

Let "S" denote the event "SME is successful".

a.


"P(S)=P(A)P(S|A)+P(B)P(S|B)"

"=0.55(0.7)+0.45(0.1)=0.43"

b.


"P(S\cap B)=P(B)P(S|B)=0.45(0.1)=0.045"

c.


"P(B|S)=\dfrac{P(B)P(S|B)}{P(A)P(S|A)+P(B)P(S|B)}"

"=\dfrac{0.045}{0.43}\approx0.10465"